A bat flying at 4.10 m/s is chasing an insect flying in the same direction. If the bat emits a 40.0-kHz chirp and receives back an echo at 40.4 kHz, what is the speed of the insect? (Take the speed of sound in air to be 340 m/s.)
_______m/s
Answer #1

Relative velocity of a bat and an insect (Rv)
f=emits frequency , fo=echo frequency.
Rv= v(fo-f)/(fo+f) [m/s] , (v=343m/s , f=40.0kHz , fo=40.4kHz)
Rv= 340(40.4-40.0)/(40.4+40.0)
Rv= 340*0.4/80.4= 1.691m/s
vi= bat speed-Rv
vi= 4.10-1.691= 3.648= 2.40[m/s] (insect away speed)